# Linear-Cost Vecchia Approximation of Multivariate Normal Probabilities

### v2

@@ 3.2 Transforming SOV and MET
Algorithm 1 describes the MC sampling stage of our transformed Vecchia MET (VMET) based on the sparse matrix $\mathbf{A}$ and the conditional standard deviation $\mathbf{l}$. The exponential tilting parameter $\bm{\gamma}$ can be obtained as discussed in Section 4 during the pre-processing stage, or it can be set as $\bm{\gamma}=\mathbf{0}$ to obtain Vecchia SOV (VSOV) as a special case.
@@
3:     if $i>1$ then
4:        $\mu_{i}\leftarrow\mathbf{A}_{i,:}\mathbf{x}$
5:     else
6:        $\mu_{i}\leftarrow 0$
7:     end if
8:     $\tilde{a}_{i}\leftarrow\frac{a_{i}-\mu_{i}}{l_{i}}-\gamma_{i}$, $\tilde{b}_{i}\leftarrow\frac{b_{i}-\mu_{i}}{l_{i}}-\gamma_{i}$, $p_{i}\leftarrow\Phi(\tilde{b}_{i})-\Phi(\tilde{a}_{i})$
9:     $y_{i}\leftarrow\Phi^{-1}(w_{i}\cdot p_{i}+\Phi(\tilde{a}_{i}))+\gamma_{i}$, $x_{i}\leftarrow\mu_{i}+y_{i}\cdot l_{i}$
10:     $h\leftarrow h\cdot p_{i}\cdot\phi(y_{i})/\phi(y_{i}-\gamma_{i})$
11:  end for
12:  if Sampling from TMVN then
13:     return  $h$ and $\mathbf{x}$
14:  else
15:     return  $h$
16:  end if
@@
Here, we use $\mathbf{w}$ to denote a sample from the unit hypercube in $\mathbb{R}^{n}$, offering users the flexibility of choosing different MC rules. Compared to MET, the complexity of each iteration (i.e., for each $i$) is reduced from $\mathcal{O}(n)$ to $\mathcal{O}(m)$ employing the sparsity of $\mathbf{A}$ in Step 4. Therefore, the per-sample complexity is reduced from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ during the MC sampling stage. We refer to Algorithm 1 as the ‘integrand’ to distinguish from the ‘method’, which includes both pre-processing and MC sampling. Algorithm 1 can be also used for accept-reject sampling of TMVN distributions, where $\mathbf{x}$ is a sample generated from the proposal density $q(\mathbf{x})$ in (9) and $h$ is the PDF ratio in (10). VMET allows us to fully exploit the sparseness of $\mathbf{V}$ by using the original integration variable $\mathbf{x}$, whereas MET uses the transformed $\mathbf{y}=\mathbf{L}^{-1}\mathbf{x}$. Since the Vecchia approximation amounts to a valid MVN density with (slightly) different covariance, Algorithm 1 enjoys the same analytical properties and efficacy in the variance reduction of the integrand as the original MET method (Botev, 2017). This is supported by our theoretical and numerical results in Sections 4 and 6, respectively.
3.3 Extension to multivariate Student-$t$ (MVT) probabilities
MVT probabilities can be defined as scale mixtures of MVN probabilities, $T_{n}(\mathbf{a},\mathbf{b};\bm{\Sigma},\nu)=E_{r}[\Phi_{n}(r\frac{\mathbf{a}}{\sqrt{\nu}},r\frac{\mathbf{b}}{\sqrt{\nu}};\bm{\Sigma})]$, where $T_{n}(\mathbf{a},\mathbf{b};\bm{\Sigma},\nu)$ denotes an MVT probability and $r$ is the square root of a $\chi^{2}$ random variable with $\nu$ degrees of freedom. The MET proposal density for MVN probabilities can be modified for multivariate Student-$t$ (MVT) probabilities by prepending the proposal density in (9) with another univariate truncated normal. Specifically, the proposal density and the likelihood ratio become:
$q_{t}(\mathbf{y};\bm{\gamma})={}$ $\frac{\mathbbm{1}_{r\in(0,\infty)}}{1-\Phi(-\gamma_{0})}\phi(r-\gamma_{0})\prod_{i=1}^{n}\frac{\mathbbm{1}_{y_{i}\in(r\tilde{a}_{i},r\tilde{b}_{i})}}{\Phi(r\tilde{b}_{i}-\gamma_{i})-\Phi(r\tilde{a}_{i}-\gamma_{i})}\phi(y_{i}-\gamma_{i}),$ (19)
$h_{t}(\mathbf{y};\bm{\gamma})={}$ $\frac{2^{1-\nu/2}(1-\Phi(-\gamma_{0}))}{\Gamma(\nu/2)}\frac{r^{\nu-1}e^{-r^{2}/2}}{\phi(r-\gamma_{0})}$ (20)
$\prod_{i=1}^{n}\left\{\Phi(r\tilde{b}_{i}-\gamma_{i})-\Phi(r\tilde{a}_{i}-\gamma_{i})\right\}\frac{\phi(y_{i})}{\phi(y_{i}-\gamma_{i})}\mathbbm{1}_{y_{i}\in(r\tilde{a}_{i},r\tilde{b}_{i})},$ (21)
respectively, and $\gamma_{0}$ is the exponential tilting parameter for $r$. These expressions were first proposed in Botev & L’Ecuyer (2015) and are included here using our Vecchia approach and notation. A numerical study comparing existing methods for estimating MVT probabilities, similar to that in Figure 3 for MVN probabilities, is shown in Figure LABEL:fig:low_dim_exp_T in the Appendix.
@@ 4 Solving For Exponential Tilting Parameters
$\hat{\bm{\gamma}}=\operatorname*{arg\,min}\max_{\mathbf{a}\leq\mathbf{x}\leq\mathbf{b}}\psi(\mathbf{x};\bm{\gamma})$ (22)
@@
The integrand $h(\mathbf{x};\hat{\bm{\gamma}})=\exp\psi(\mathbf{x};\hat{\bm{\gamma}})$ has vanishing relative error. That is
@@ 5.1 Variable Reordering Based on Vecchia
An integration-variable reordering method, referred to as univariate variable reordering, based on the correlation structure in $\bm{\Sigma}$ and the integration limits $\mathbf{a}$ and $\mathbf{b}$, was described in Gibson et al. (1994) to empirically improve the MC sampling variance. Univariate reordering, which is provided in Algorithm 2, iteratively computes conditional univariate normal probabilities and prioritizes the one with the smallest probability, resulting in $\mathcal{O}(n^{3})$ complexity.
Input: $n,\mathbf{a},\mathbf{b},\bm{\Sigma}$
Result: New variable order $\mathbf{r}$
1:  Initialize $\mathbf{r}\leftarrow[1,\ldots,n]$, $\bm{\sigma}\leftarrow diag(\bm{\Sigma})^{1/2}$, $\bm{\mu}\leftarrow\mathbf{0}_{n}$
2:  for $i=1,2,\ldots,n$ do
3:     $j\leftarrow\operatorname*{arg\,min}_{j\geq i}\Phi((b_{j}-\mu_{j})/\sigma_{j})-\Phi((a_{j}-\mu_{j})/\sigma_{j})$
4:     Switch the $i$-th and $j$-th coefficient in $\mathbf{a}$, $\mathbf{b}$, $\bm{\sigma}$, $\bm{\mu}$, and $\mathbf{r}$
5:     Compute $E_{X\sim N(\mu_{i},\sigma_{i})}[X\mid a_{i}<X<b_{i}]$ and assign to $\mu_{i}$
6:     Under $\mathbf{x}\sim N_{n}(\mathbf{0},\bm{\Sigma})$, compute $\mu_{j}\leftarrow E[x_{j}\mid\mathbf{x}_{1:i}=\bm{\mu}_{1:i}]$
$\sigma_{j}\leftarrow\sqrt{\mathrm{Var}[x_{j}\mid\mathbf{x}_{1:i}=\bm{\mu}_{1:i}]}$, for $j=i+1,\ldots,n$
7:  end for
8:  return  $\mathbf{r}$
Algorithm 2 Univariate variable reordering (Gibson et al., 1994)
We propose a fast Vecchia-based version of the univariate variable reordering in Algorithm 3.
Input: $n,\mathbf{a},\mathbf{b},\bm{\Sigma}$
Result: New variable order $\mathbf{r}$
1:  Standardize $\mathbf{a},\mathbf{b},\bm{\Sigma}$ s.t. $diag(\bm{\Sigma})=\mathbf{1}$
2:  Initialize $\mathbf{r}\leftarrow[1,\ldots,n]$, $\bm{\sigma}\leftarrow diag(\bm{\Sigma})^{1/2}$, $\bm{\mu}\leftarrow\mathbf{0}_{n}$, $c(j)\leftarrow\{\}$ for $j=1,\ldots,n$
3:  for $i=1,2,\ldots,n$ do
4:     $j\leftarrow\operatorname*{arg\,min}_{j\geq i}\Phi((b_{j}-\mu_{j})/\sigma_{j})-\Phi((a_{j}-\mu_{j})/\sigma_{j})$
5:     Switch the $i$-th and $j$-th coefficient in $\mathbf{a}$, $\mathbf{b}$, $\bm{\sigma}$, $\bm{\mu}$, and $\mathbf{r}$
6:     Compute $E[X\mid X\sim N(\mu_{i},\sigma_{i}),a_{i}<X<b_{i}]$ and assign to $\mu_{i}$
7:     for $j=i+1,\ldots,n$ do
8:        if $|c(j)|<m$ or $\min(\bm{\Sigma}_{s_{j},\mathbf{s}_{c(j)}})<\bm{\Sigma}_{s_{j},s_{i}}$ then
9:           $c(j)\leftarrow c(j)\bigcup\{i\}$
10:           if $|c(j)|>m$ then
11:              $k\leftarrow\operatorname*{arg\,min}\bm{\Sigma}_{s_{j},\mathbf{s}_{c(j)}}$, remove the $k$-th coefficient in $c(j)$
12:           end if
13:           $\mu_{j}\leftarrow E[x_{j}\mid\mathbf{x}_{c(j)}=\bm{\mu}_{c(j)}]$, $\sigma_{j}\leftarrow\sqrt{\mathrm{Var}[x_{j}\mid\mathbf{x}_{c(j)}=\bm{\mu}_{c(j)}]}$
14:        end if
15:     end for
16:  end for
17:  return  $\mathbf{r}$
Algorithm 3 Vecchia-based variable reordering
@@
Under the Vecchia approximation, the conditional expectation and variance in Line 2 of Algorithm 2 can be computed in $\mathcal{O}(m^{3})$ time for each $j$ as indicated by Line 13 of Algorithm 3, amounting to a total algorithm complexity of $\mathcal{O}(n^{2})$. Because the conditioning sets in the Vecchia approximation are subsets of previous indices (i.e., $c(i)\subset\{1,\ldots,i-1\}$), we need to construct and update the conditioning sets during univariate reordering as described in Lines 9 and 11 of Algorithm 3, where the conditioning sets are chosen as the $m$ previously indexed integration variables that have the strongest correlation with the $j$-th integration variable.
@@
Variable reordering can be reduced to linear complexity under a special case of the Vecchia approximation with $c(i)=\{1,\ldots,m\}$ for $i>m$, which results in the fully independent conditional (FIC) approach (Snelson & Ghahramani, 2007; Katzfuss & Guinness, 2021), one of the most popular GP approximations in machine learning. Under the FIC assumption, the variable ordering (after selecting the first $m$ variables) can be obtained in a single iteration, hence achieving an overall $\mathcal{O}(n)$ complexity. For the first $m$ variables, we can simply run the loop in Algorithm 3 for $m$ iterations. We refer to this $\mathcal{O}(n)$ ordering as the FIC-based variable reordering to distinguish it from the $\mathcal{O}(n^{2})$ Vecchia-based variable reordering in Algorithm 3. We provide a numerical comparison on the classic univariate reordering, the Vecchia-based variable reordering, and the FIC-based variable reordering in Section 6.1.
@@ 5.2 Multi-level Monte Carlo sampling
where $h_{l}^{m_{1}}$ and $h_{l}^{m_{2}}$ are correlated for each $l=1,\ldots,N_{2}$ by using the same randomness profile to reduce the impact from the variance of $h$. The resulting multi-level-MC probability estimate is $\hat{h}=\frac{1}{N_{1}}\sum_{l=1}^{N_{1}}h_{l}^{m_{1}}-\hat{\epsilon}$. The dependence between $\frac{1}{N_{1}}\sum_{l=1}^{N_{1}}h_{l}^{m_{1}}$ and $\hat{\epsilon}$ is minimal, because $N_{2}\ll N_{1}$. Multi-level MC heuristically improves the accuracy of the VMET method, but its practical performance may depend on the distribution of $h_{l}^{m_{1}}-h_{l}^{m_{2}}$. In certain cases, the bias of the VMET integrand may not monotonically decrease with $m$, which was observed in our numerical simulations (e.g., see Figures 3 and 4). This may render multi-level MC less effective in the bias reduction for VMET integrands, because high-cost samples may not result in smaller bias and hence, we categorize multi-level MC as an optional technique and demonstrate its impact under a constant correlation structure in Section 6.4.
@@ 6.1 Variable reordering comparison
In this section, we show that the variable reordering heuristic proposed for SOV (Gibson et al., 1994) can also improve the results for MET and VMET. We demonstrate that our proposed Vecchia-based variable reordering (Algorithm 3) can achieve the same level of improvement in reducing the MC variance as the univariate reordering (Algorithm 2). SOV and TLR are not considered in this section, because their benefits from variable reordering have been documented in the literature (e.g., Genz & Bretz, 2009; Cao et al., 2022).
@@
Figure 1 compares the MET and VMET (with $m=30$) integrands with and without univariate variable reordering, where each boxplot consists of thirty estimates and each estimate used $N=10{,}000$ MC samples.
Figure 1: Impact of variable reordering on MET and VMET integrands for the three scenarios described in Table 1. Each boxplot consists of $30$ estimates of the same MVN probability. Each panel compares MET and VMET integrands with and without univariate reordering.
Variable reordering reduced the variability of both integrands, hence improving the accuracy for MVN probability estimation. The effectiveness of variable reordering varied across different MVN probabilities. In Scenario 3, where integration limits were centered and the same across integration variables, variable reordering had a smaller impact than in the other two cases. Furthermore, our proposed VMET integrand had overall negligible bias and the same level of variability compared with the MET integrand.
@@ 6.2 Comparison for low-dimensional MVN probabilities
Figure 3: For the scenarios in Table 1 with $n=900$, we show the RMSE of 30 probability estimates against the computation time (both on a log scale), for VMET, VCDF, TLR, SOV, and MET. VMET is run with $m\in\{10,30,50,70,90\}$ and $N=10^{4}$, SOV and TLR are run with $N\in\{1,3,5,7,9\}\times 10^{4}$, and MET and VCDF are run with $N\in\{6,7,8,9,10\}\times 10^{3}$. The average of MET estimates with $N=10^{4}$ was used as the truth.
@@ 6.3 Comparison for high-dimensional MVN probabilities
For larger $n$, we only compared our method to TLR and SOV, as MET and VCDF are computationally infeasible. We considered three $n=6{,}400$-dimensional MVN probabilities according to Scenarios 1 to 3 in Table 1, respectively. For this high-dimensional study, the nugget parameter was increased from $0.01$ to $0.03$ to avoid singularity, while the other covariance parameters remained unchanged, and we considered $m$ between $30$ and $70$ for VMET. The RMSE of log-probability estimates and computation times are shown in Figure 4.
@@
For Scenarios 1 and 2, we considered VMET with $m=70$ as the benchmark, as VMET arguably has lower MC variance than SOV and TLR for tail MVN probabilities, and the bias caused by the Vecchia approximation was likely insignificant compared with the MC variance. For Scenario 3, we used SOV as benchmark, because for centered MVN probabilities, MET and SOV amount to the same integrand, hence SOV is expected to have the same accuracy level as MET. Compared to SOV and TLR, VMET significantly reduced the MC variance, hence improving estimation accuracy for tail MVN probabilities; for centered MVN probabilities, VMET also achieved a more favorable trade-off between estimation error and computation cost. Figure LABEL:fig:high_dim_exp_boxplot in the Appendix provides more detailed information on the three methods’ probability estimates. Furthermore, TLR, the state-of-the-art in scalability, may suffer from unpredictable numerical singularity, whereas VMET is numerically stable, because unlike low-rank matrix approximations, the Vecchia approximation has a theoretical guarantee for positive definiteness.

### v3

@@ 3.2 Transforming SOV and MET
With Proposition 2, we transform the SOV and MET algorithms such that MC sampling can be performed at linear complexity as described in Algorithm 1. The new method is coined Vecchia MET (VMET). The exponential tilting parameter $\bm{\gamma}$ can be obtained as discussed in Section 4 during the pre-processing stage, or it can be set as $\bm{\gamma}=\mathbf{0}$ to obtain Vecchia SOV (VSOV) as a special case. VMET uses the original integration variable $\mathbf{x}$, and hence we define $q(\mathbf{x})=q^{B}(\mathbf{L}^{-1}\mathbf{x})$ and $h(\mathbf{x})=h^{B}(\mathbf{L}^{-1}\mathbf{x})$ as the proposal density and the density ratio, up to a normalizing constant, used in VMET.
@@
3:     if $i>1$, $\mu_{i}\leftarrow\mathbf{A}_{i,:}\mathbf{x}$  else $\mu_{i}\leftarrow 0$
4:     $\tilde{a}_{i}\leftarrow\frac{a_{i}-\mu_{i}}{l_{i}}-\gamma_{i}$, $\tilde{b}_{i}\leftarrow\frac{b_{i}-\mu_{i}}{l_{i}}-\gamma_{i}$, $p_{i}\leftarrow\Phi(\tilde{b}_{i})-\Phi(\tilde{a}_{i})$
5:     $y_{i}\leftarrow\Phi^{-1}(w_{i}\cdot p_{i}+\Phi(\tilde{a}_{i}))+\gamma_{i}$, $x_{i}\leftarrow\mu_{i}+y_{i}\cdot l_{i}$
6:     $h\leftarrow h\cdot p_{i}\cdot\phi(y_{i})/\phi(y_{i}-\gamma_{i})$
7:  end for
8:  if Sampling from TMVN then
9:     return  $h$ and $\mathbf{x}$
10:  else
11:     return  $h$
12:  end if
@@
We use $\mathbf{w}$ to denote a sample from the unit hypercube in $\mathbb{R}^{n}$, offering the flexibility of choosing different MC rules. Compared to MET, the complexity of each iteration (i.e., for each $i$) is reduced from $\mathcal{O}(n)$ to $\mathcal{O}(m)$ employing the sparsity of $\mathbf{A}$ in Step 3. Therefore, the per-sample complexity is reduced from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ during the MC sampling stage. We refer to Algorithm 1 as the ‘integrand’ to distinguish from the ‘method’, which includes both pre-processing and MC sampling. Algorithm 1 can be also used for accept-reject sampling of TMVN distributions, where $\mathbf{x}$ is a sample generated from the proposal density $q(\mathbf{x})$ and $h$ is the PDF ratio. VMET allows us to fully exploit the sparseness of $\mathbf{V}$ by using the original integration variable $\mathbf{x}$. Since the Vecchia approximation amounts to a valid MVN density with (slightly) different covariance, Algorithm 1 enjoys the same analytical properties and efficacy in the variance reduction of the integrand as the original MET method (Botev, 2017). This is supported by our theoretical and numerical results in Sections 4 and 6, respectively. MVT probabilities can be defined as scale mixtures of MVN probabilities. Based on this, we provide modifications of the proposed VMET integrand for MVT probabilities, along with associated numerical results, in Appendix LABEL:sec:appendix_MVT.
@@ 4 Solving For Exponential Tilting Parameters
$\hat{\bm{\gamma}}=\operatorname*{arg\,min}\max_{\mathbf{a}\leq\mathbf{x}\leq\mathbf{b}}\psi(\mathbf{x};\bm{\gamma})$
@@
The integrand $h(\mathbf{x};\hat{\bm{\gamma}})=\exp\psi(\mathbf{x};\hat{\bm{\gamma}})$ has vanishing relative error:
@@ 5.1 Variable Reordering Based on Vecchia
An integration-variable reordering method, referred to as univariate variable reordering was described in Gibson et al. (1994) to empirically improve the MC sampling variance. Univariate reordering, provided in Algorithm LABEL:alg:univar_reorder in the appendix for completeness, iteratively computes conditional univariate normal probabilities and prioritizes the one with the smallest probability, resulting in $\mathcal{O}(n^{3})$ complexity. We propose a fast Vecchia-based version of the univariate variable reordering in Algorithm LABEL:alg:vecc_reorder. Under the Vecchia approximation, the conditional expectation and variance in Line LABEL:stp:cond_mean_var of Algorithm LABEL:alg:univar_reorder can be computed in $\mathcal{O}(m^{3})$ time for each $j$ as indicated by Line LABEL:stp:cond_mean_var_vecc of Algorithm LABEL:alg:vecc_reorder, amounting to a total algorithm complexity of $\mathcal{O}(n^{2})$. Because the conditioning sets in the Vecchia approximation are subsets of previous indices (i.e., $c(i)\subset\{1,\ldots,i-1\}$), we need to construct and update the conditioning sets during univariate reordering as described in Lines LABEL:stp:cond_set_update_1 and LABEL:stp:cond_set_update_2 of Algorithm LABEL:alg:vecc_reorder, where the conditioning sets are chosen as the $m$ previously indexed integration variables that have the strongest correlation with the $j$-th integration variable.
@@
Variable reordering can be reduced to linear complexity under a special case of the Vecchia approximation with $c(i)=\{1,\ldots,m\}$ for $i>m$, which results in the fully independent conditional (FIC) approach (Snelson & Ghahramani, 2007; Katzfuss & Guinness, 2021), a popular GP approximation in machine learning. Under the FIC assumption, the variable ordering (after selecting the first $m$ variables) can be obtained in a single iteration, hence achieving an overall $\mathcal{O}(n)$ complexity. For the first $m$ variables, we can simply run the loop in Algorithm LABEL:alg:vecc_reorder for $m$ iterations. We refer to this $\mathcal{O}(n)$ ordering as the FIC-based variable reordering to distinguish it from the $\mathcal{O}(n^{2})$ Vecchia-based variable reordering in Algorithm LABEL:alg:vecc_reorder. We provide a numerical comparison of the univariate reordering, the Vecchia-based variable reordering, and the FIC-based variable reordering in Section 6.1.
@@ 5.2 Multi-level Monte Carlo sampling
where $h_{l}^{m_{1}}$ and $h_{l}^{m_{2}}$ are correlated for each $l=1,\ldots,N_{2}$ by using the same randomness profile to reduce the impact from the variance of $h$. The resulting multi-level-MC probability estimate is $\hat{h}=\frac{1}{N_{1}}\sum_{l=1}^{N_{1}}h_{l}^{m_{1}}-\hat{\epsilon}$. The dependence between $\frac{1}{N_{1}}\sum_{l=1}^{N_{1}}h_{l}^{m_{1}}$ and $\hat{\epsilon}$ is minimal, because $N_{2}\ll N_{1}$. Multi-level MC heuristically improves the accuracy of the VMET method, but its practical performance may depend on the distribution of $h_{l}^{m_{1}}-h_{l}^{m_{2}}$. In certain cases, the bias of the VMET integrand may not monotonically decrease with $m$, which was observed in our numerical simulations (e.g., see Figures 2 and 3). This may render multi-level MC less effective in the bias reduction, because high-cost samples may not result in smaller bias and hence, we categorize multi-level MC as an optional technique and demonstrate its impact under a constant correlation structure in Section 6.4.
@@ 6.1 Variable reordering comparison
In this section, we show that the variable reordering heuristic proposed for SOV (Gibson et al., 1994) can also improve the results for MET and VMET. We demonstrate that our proposed Vecchia-based variable reordering (Algorithm LABEL:alg:vecc_reorder) can achieve the same level of improvement in reducing the MC variance as the univariate reordering (Algorithm LABEL:alg:univar_reorder). SOV and TLR are not considered in this section, because their benefits from variable reordering have been documented in the literature (e.g., Genz & Bretz, 2009; Cao et al., 2022).
@@
Figure LABEL:fig:order_cmp_MET_VMET in the appendix compares the MET and VMET (with $m=30$) integrands with and without univariate variable reordering. Variable reordering reduced the variability of both integrands, hence improving the accuracy for MVN probability estimation. The effectiveness of variable reordering varied across different MVN probabilities. In Scenario 3, where integration limits were centered and the same, variable reordering had a smaller impact than in the other two cases. Furthermore, our proposed VMET integrand had overall negligible bias and the same level of variability compared with the MET integrand.
@@ 6.2 Comparison for low-dimensional MVN probabilities
Figure 2: For the scenarios in Table 1 with $n=900$, we show the RMSE of 30 probability estimates against the computation time (both on a log scale), for VMET, VCDF, TLR, SOV, and MET. VMET is run with $m\in\{10,30,50,70,90\}$ and $N=10^{4}$, SOV and TLR are run with $N\in\{1,3,5,7,9\}\times 10^{4}$, and MET and VCDF are run with $N\in\{6,7,8,9,10\}\times 10^{3}$. An MET estimate with $N=10^{7}$ is used as the truth.
@@ 6.3 Comparison for high-dimensional MVN probabilities
For larger $n$, we only compared our method to TLR and SOV, as MET and VCDF are computationally infeasible. We considered three $n=6{,}400$-dimensional MVN probabilities according to Scenarios 1 to 3 in Table 1, respectively. For this high-dimensional study, the nugget parameter was increased from $0.01$ to $0.03$ to avoid singularity, while the other covariance parameters remained unchanged. The RMSE of log-probability estimates and computation times are shown in Figure 3.
@@
For Scenarios 1 and 2, we considered VMET with $m=70$ as the benchmark, as VMET arguably has lower MC variance than SOV and TLR for tail MVN probabilities, and the bias caused by the Vecchia approximation was likely insignificant compared with the MC variance. For Scenario 3, we used SOV as benchmark, because for centered MVN probabilities, MET and SOV amount to the same integrand, hence SOV is expected to have the same accuracy as MET. Compared to SOV and TLR, VMET significantly reduced the MC variance, hence improving estimation accuracy for tail MVN probabilities; for centered MVN probabilities, VMET also achieved a more favorable trade-off between estimation error and computation cost. Figure LABEL:fig:high_dim_exp_boxplot in the Appendix provides more detailed information on the three methods’ probability estimates. Furthermore, TLR, the state-of-the-art in scalability, may suffer from unpredictable numerical singularity, whereas VMET is numerically stable, because unlike low-rank matrix approximations, the Vecchia approximation has a theoretical guarantee for positive definiteness (see Prop. 1 in Katzfuss & Guinness, 2021).